# Euler systems and relative Satake isomorphism

### v1

@@ Abstract.
We explain how the tame norm relation in Euler systems is a formal consequence of the unramified Plancherel formula in the relative Langlands program and a mild refinement of the generalized Cartan decomposition. This uniformly recovers many of the known Euler systems and produces a new split anticyclotomic Euler system in a case studied by Cornut.
@@ 1. Introduction
The Bloch–Kato conjecture suggests a deep connection between the arithmetic of motives and special values of $L$-functions. The method of Euler systems is a method of understanding this connection. An Euler system consists of a family of motivic classes satisfying two types of relations: “wild” ones for relations in a $p$-adic tower, and “tame” ones for relations at places away from $p$. Experiences suggest that the wild norm relations require quite special conditions on $p$, such as some form of ordinarity. On the other hand, tame norm relations seem to exist in much greater generalities, and they are already enough to deduce some cases of the Bloch–Kato conjecture, cf. [JNS24, LS24]. However, in the many known examples of Euler systems, the construction of classes satisfying the tame norm relations usually requires some ad hoc choices followed by extensive case-by-case calculations.
@@
$\mathrm{U}(n)\times\mathrm{U}(n+1)$ $\mathrm{U}(n)\backslash\mathrm{U}(n)\times\mathrm{U}(n+1)$ Lai–Skinner [LS24]
$\mathrm{U}(2n)$ $\mathrm{U}(n)\times\mathrm{U}(n)\backslash\mathrm{U}(2n)$ Graham–Shah [GS23]
$\mathrm{SO}(2n+1)$ $\mathrm{U}(n)\backslash\mathrm{SO}(2n+1)$ Cornut [Cor18]
@@
$\operatorname{Res}_{F/\mathbb{Q}}\mathrm{GL}_{2}$uparrow Grossi [Gro20]
$\mathrm{GU}(2,1)$uparrow Loeffler–Skinner–Zerbes [LSZ22a]
@@
The idea of using spherical varieties in the construction of Euler systems goes back to Cornut [Cor18]. Unfortunately, there is a gap in his construction related to the definition of Hecke operators, as explained in [Sha23]. For this case, we produce a new split anticyclotomic Euler system and provide an independent proof of the main arithmetic application on Bloch-Kato conjecture.
@@ 1.1. Simple case of main result
We will state the application of our results for the pushforward of cycles, covering the first three rows of the above table. We construct split anticyclotomic Euler systems in the sense of Jetchev–Nekovář–Skinner [JNS24], which we will simply call JNS Euler systems.
@@
holds. After fixing level structures for $\mathbf{G}$ and $\mathbf{H}$, this defines a special cycle on $\mathrm{Sh}_{\mathbf{G}}$ in the arithmetic middle dimension, which we expect is related to the central derivative of an $L$-function.
We set up some notations for a JNS Euler system. Let $\mathscr{L}$ be the set of places of $F$ which split in $E$, with an explicit finite set of exceptions. For each $\ell\in\mathscr{L}$, fix a place $\lambda$ of $E$ above $\ell$, and let $\mathrm{Frob}_{\lambda}$ be the arithmetic Frobenius at $\lambda$. Let $\mathscr{R}$ be the set of square-free products of places in $\mathscr{L}$. For $\mathfrak{m}\in\mathscr{R}$, let $E[\mathfrak{m}]$ be the ring class field of conductor $\mathfrak{m}$, so it is associated to the order $\mathcal{O}_{F}+\mathfrak{m}\mathcal{O}_{E}$ by class field theory. Fix a minuscule representation $V$ of the dual group $\check{G}$ of $\mathbf{G}$. For each $\ell$, let $\mathcal{H}_{V,\ell}(X)$ be the Hecke polynomial of $V$ at the place $\ell$ (cf. Example 3.12).
@@
Let $d=\dim\mathrm{Sh}_{\mathbf{G}}$. Suppose $\mathbf{X}=\mathbf{H}\backslash\mathbf{G}$ is a spherical $\mathbf{G}$-variety, and there is a character $\nu:\mathbf{H}\to\mathrm{U}(1)$ satisfying the “combinatorially trivial” condition of Example 2.3 such that the induced Shimura datum on $\mathrm{U}(1)$ is non-trivial. Then there exists a collection of classes
@@
$\operatorname{Tr}_{E[\mathfrak{m}]}^{E[\mathfrak{m}\ell]}z_{\mathfrak{m}\ell}=\mathcal{H}_{V,\ell}(\mathrm{Frob}_{\lambda}^{-1})\cdot z_{\mathfrak{m}}.$
@@
We can in fact replace $\mathbb{Z}_{p}$ by more general coefficient systems. Under some automorphic assumptions, a standard Abel–Jacobi map procedure described in §4.3 gives the required Euler system in Galois cohomology.
@@
Let $\pi$ be a stable cohomological automorphic representation of $\mathbf{G}(\mathbb{A}_{F})$ distinguished by $\mathbf{X}$. Let $\rho_{\pi}$ be the $p$-adic Galois representation attached to $\pi$ and the Shimura cocharacter for $\mathbf{G}$.
@@
  • Kottwitz’s conjecture is used to show that $\rho_{\pi}$ actually contributes to the cohomology of $\mathrm{Sh}_{\mathbf{G}/\bar{E}}$. In all of our cases, the Shimura variety is of abelian type, and what we need follows from the work of Kisin–Shin–Zhu [KSZ21].
@@
We also have a version of the above results for the pushforward of Eisenstein classes (Example 4.3), which recovers the final four entries in the table. However, they depend on numerical consequences of the results of Sakellaridis–Wang [SW22] which are not yet available in the mixed characteristics setting.
@@
1.2. Idea of proof
The theorem above is a combination of the following two steps which are completely different in nature:
@@ 1.2. Idea of proof
    Construction of a “motivic theta series” (Definition 4.1).
@@
    Local harmonic analysis on spherical varieties (Proposition 3.13).
We now explain each item in turn.
1.2.1. Motivic theta series
Observation 1.6 (Definition 4.1, Propositions 4.8, 4.11).
where $\mathbf{X}$ is a spherical variety.
In the cycles case, we take $\mathbf{X}=\mathbf{H}\backslash\mathbf{G}$, so we are interested in the function space $C_{c}^{\infty}(\mathbf{H}\backslash\mathbf{G},\mathbb{Z}_{p})$. In previous works, the source space is often taken to be a space of coinvariants $C_{c}^{\infty}(\mathbf{G},\mathbb{Z}_{p})_{\mathbf{H}}$. These two spaces are equal after tensoring with $\mathbb{Q}_{p}$, but they are different integrally. In other words, our observation is that $C_{c}^{\infty}(\mathbf{X},\mathbb{Z}_{p})$ is the more natural integral structure. The cases involving Eisenstein classes can be treated similarly, taking $\mathbf{X}$ to be the vector bundle
This observation was already used by the third named author in [LS24] to bypass some Iwasawa-theoretic arguments needed previously to deal with torsions.
1.2.2. Local harmonic analysis
Once such a motivic theta series exist, the problem of tame norm relations is entirely reduced to constructing test vectors satisfying norm relations in $C_{c}^{\infty}(X,\mathbb{Z}_{p})$, where $X=\mathbf{X}(F_{\ell})$.
(†) is it true that $\text{is it true that }\mathcal{H}_{\ell}\cdot\Phi_{0}\in\operatorname{Tr}_{K^{0}}^{K^{1}}C_{c}^{\infty}(X,\mathbb{Z}_{p})^{K^{1}}?$
This is stronger than just requiring the function $\mathcal{H}_{\ell}\cdot\Phi_{0}$ to take values in $\mathbb{Z}_{p}$. Indeed, in the extreme case where the $K^{0}$-orbit of a point $x\in X$ coincides with its $K^{1}$-orbit, we need the value at $x$ to be divisible by the index $[K^{0}:K^{1}]=\ell-1$. We give a necessary condition for when this kind of additional divisibility requirements can occur, in terms of the geometry of $\mathbf{X}$ (Proposition 3.4). This is done by extending the proof of the generalized Cartan decomposition due to Gaitsgory–Nadler [GN10, Theorem 8.2.9] and Sakellaridis [Sak12, Theorem 2.3.8],
Observation 1.7 (Proposition 3.9).
In the cases considered in this paper, the formula for the inverse relative Satake isomorphism gives the values of the function $\mathcal{H}_{\ell}\cdot\Phi_{0}$ as polynomials in $\ell$. Now view $\ell$ as a formal variable. Instead of computing this polynomial, we specialize at $\ell=1$ and show by an anti-symmetry argument that the result is 0 at certain points in $X$ (along “walls of type T”). This proves its divisibility by $\ell-1$ at those points. This additional divisibility is a new phenomenon in the relative setting and not present in the classical Satake isomorphism (Remark 3.10).

### v2

@@ Abstract.
We explain how the unramified Plancherel formula in the relative Langlands program gives a natural way of constructing test vectors which satisfy the tame norm relations of an Euler system. This uniformly recovers many of the known Euler systems, and in the twisted Friedberg–Jacquet setting, we produce a new split anticyclotomic Euler system.
@@ 1. Introduction
The Bloch–Kato conjecture suggests a deep connection between the arithmetic of motives and special values of $L$-functions. The method of Euler systems is one way of understanding this connection. An Euler system consists of a family of motivic classes satisfying two types of relations: “wild” ones for relations in a $p$-adic tower, and “tame” ones for relations at places away from $p$. Experiences suggest that the wild norm relations require quite special conditions on $p$, such as some form of ordinarity. On the other hand, tame norm relations seem to exist in much greater generalities, and they are already enough to deduce some cases of the Bloch–Kato conjecture, cf. [JNS24, LS24]. However, in the many known examples of Euler systems, the construction of classes satisfying the tame norm relations usually requires some ad hoc choices followed by extensive case-by-case calculations.
@@
[…]
Inner form of $\uparrow$ Twisted Friedberg–Jacquet
@@
$\operatorname{Res}_{F/\mathbb{Q}}\mathrm{GL}_{2}$ $\uparrow$ Grossi [Gro20]
$\mathrm{GU}(2,1)$ $\uparrow$ Loeffler–Skinner–Zerbes [LSZ22a]
@@ 1.1. Simple case of main result
We will state the application of our results for the pushforward of cycles, covering the first three rows of Table 1. We construct split anticyclotomic Euler systems in the sense of Jetchev–Nekovář–Skinner [JNS24], which we will simply call JNS Euler systems.
@@
holds. After fixing level structures for $\mathbf{G}$ and $\mathbf{H}$, this defines a special cycle on $\mathrm{Sh}_{\mathbf{G}}$ in the arithmetic middle dimension, which is expected to be related to the central derivative of an $L$-function.
We set up some notations for a JNS Euler system. Let $\mathscr{L}$ be the set of places of $F$ which split in $E$, with an explicit finite set of exceptions. For each $\ell\in\mathscr{L}$, fix a place $\lambda$ of $E$ above $\ell$, and let $\mathrm{Frob}_{\lambda}$ be the arithmetic Frobenius at $\lambda$. Let $\mathscr{R}$ be the set of square-free products of places in $\mathscr{L}$. For $\mathfrak{m}\in\mathscr{R}$, let $E[\mathfrak{m}]$ be the ring class field of conductor $\mathfrak{m}$, so it is associated to the order $\mathcal{O}_{F}+\mathfrak{m}\mathcal{O}_{E}$ by class field theory. Our first result is the construction of a motivic Euler system.
@@
Let $d=\dim\mathrm{Sh}_{\mathbf{G}}$. Suppose $\mathbf{X}=\mathbf{H}\backslash\mathbf{G}$ is a spherical $\mathbf{G}$-variety. Let $\nu:\mathbf{H}\to\mathrm{U}(1)$ be a character satisfying the following conditions.
  1. (1)
    The character $\nu$ is combinatorially trivial (Definition 2.2, also cf. Example 2.3).
  2. (2)
    The induced Shimura datum on $\mathrm{U}(1)$ is non-trivial.
Let $\mathcal{H}_{\ell}$ be the Hecke polynomial attached to $\mathbf{X}$ in the sense of Definition 3.17. Then there exists a collection of classes
@@
$\operatorname{Tr}_{E[\mathfrak{m}]}^{E[\mathfrak{m}\ell]}z_{\mathfrak{m}\ell}=\mathcal{H}_{\ell}(\mathrm{Frob}_{\lambda}^{-1})\cdot z_{\mathfrak{m}}.$
@@
The classes $z_{\mathfrak{m}}$ come from Hecke translates of special cycles, as usual. The novelty here is that the existence of such Hecke translates is abstractly deduced by exploiting local harmonic analysis on spherical varieties.
Remark 1.2.
We informally explain some of the definitions occurring in the above theorem.
  • In the relative Langlands program, the period integral attached to $\mathbf{X}$ is conjecturally related to an $L$-function for automorphic representations of $\mathbf{G}$. The Hecke operator $\mathcal{H}_{\ell}$ gives the local factors of this $L$-function under the classical Satake isomorphism. In all cases considered in this paper, this is the “standard” $L$-function in that setting.
    It is important to note that this is different from the “$L$-function of XX” defined in [Sak13, Definition 7.2.3]. In its notation, we only keep (a part of) the denominator and disregard the adjoint $L$-functions.
  • The conditions on $\nu$ allow us to interpret cohomology classes for $\mathrm{Sh}_{\mathbf{G}}$ over $E[\mathfrak{m}]$ as classes for $\mathrm{Sh}_{\mathbf{G}\times\mathrm{U}(1)}$ over the base field $E$. The combinatorially trivial condition guarantees that the $L$-values of all such character twists are described using the same period integral.
    In the Friedberg–Jacquet case ($\mathrm{U}(n)\times\mathrm{U}(n)\hookrightarrow\mathrm{U}(2n)$), these conditions explain the appearance of the character $\frac{\det h_{1}}{\det h_{2}}$. However, in the triple product case, the Ichino formula only describes the central value, and such a character does not exist. This is a well-known obstacle for extending the Gross–Kudla–Schoen cycle to an Euler system.
We can in fact replace $\mathbb{Z}_{p}$ by more general coefficient systems. We also have a version of the above theorem for the pushforward of Eisenstein classes (Example 4.3), which recovers the final four entries in the table. However, they depend on numerical consequences of the results of Sakellaridis–Wang [SW22], which are not yet available in the mixed characteristics setting.
1.2. Example: Twisted Friedberg–Jacquet setting
We now describe the third row of Table 1 in details and highlight some new features of our approach. In this setting, our Euler system is new.
We use the set-up in [LXZ25]. Let $F_{0}$ be a totally real field. Let $E_{0}/F_{0}$ be a totally real quadratic extension, and let $F/F_{0}$ be a CM extension. Let $E=E_{0}\otimes_{F_{0}}F$, so it is a biquadratic extension of $F_{0}$. Let $F^{\prime}/F_{0}$ be the remaining quadratic extension of $F_{0}$ contained in $E$. Let $\sigma$ denote the non-trivial element of $\operatorname{Gal}(F/F_{0})$. In summary, we have the field diagram
[…]
Let $(B,\ast)$ be a division algebra over $F$ of dimension $(2n)^{2}$ with an involution $\ast$ of second kind. This forces $B$ to be a matrix algebra at the archimedean places. Suppose that $(B,\ast)$ has signature (1,2n1)(1,2n-1) at one archimedean place and $(0,2n)$ at the other archimedean places. Suppose further that there is an embedding $E\hookrightarrow B$ whose image is fixed by *, then the centralizer $B^{E}$ is a division algebra over $E$ of dimension $n^{2}$.
Define the following algebraic groups over $F_{0}$
$\mathbf{H}=\operatorname{Res}_{E_{0}/F_{0}}\mathrm{U}(B^{E})=\{g\in(B^{E})^{\times}|g^{*}g=1\},\quad\mathbf{G}=\mathrm{U}(B)=\{g\in B^{\times}|g^{*}g=1\}$
They both have Shimura varieties with reflex fields (contained in) $F$ and dimensions
$\dim_{F}\mathrm{Sh}_{\mathbf{H}}=n-1,\quad\dim_{F}\mathrm{Sh}_{\mathbf{G}}=2n-1.$
Therefore, the basic dimension numerology is satisfied.
Let $\pi$ be a symplectic automorphic representation of $\mathrm{GL}_{2n}(\mathbb{A}_{F_{0}})$ whose base change to $F$ is cuspidal. Let $\rho_{\pi}$ be the $2n$ dimensional Galois representation attached to $\pi$. Our main result in this case constructs the tame part of a JNS Euler system for the CM extension $F^{\prime}/F_{0}$.
Theorem 1.3.
There is a JNS Euler system for the decomposable representation
$(\rho_{\pi}\oplus\rho_{\pi}\eta_{F/F_{0}})|_{\operatorname{Gal}_{F^{\prime}}}$
whose base class is the $p$-adic étale realization of the special cycle corresponding to the embedding $\mathrm{Sh}_{\mathbf{H}}\hookrightarrow\mathrm{Sh}_{\mathbf{G}}$.
This theorem is a standard consequence of Theorem 1.1, and we now explain the conditions of that theorem in this setting. Since our Euler system is for $F^{\prime}/F_{0}$, the augmentation torus is $\mathbf{T}=\mathrm{U}(1)_{F^{\prime}/F_{0}}$. The reduced norm defines a character $\mathrm{Nrd}:\mathrm{U}(B^{E})\to\mathrm{U}(1)_{E/E_{0}}$ over $E_{0}$. Using it, we obtain a character
$\nu:\mathbf{H}\to(\operatorname{Res}_{E_{0}/F_{0}}\mathrm{U}(1)_{E/E_{0}})/\mathrm{U}(1)_{F/F_{0}}\simeq\mathbf{T},$
where the final map is the norm map from $E$ to $F^{\prime}$. Using it, we define the augmented pair of groups
$\mathbf{H}\hookrightarrow\widetilde{\mathbf{G}}=\mathbf{G}\times\mathbf{T},\quad h\mapsto(h,\nu(h))$
It is easy to verify using Example 2.3 that the character $\nu$ is combinatorially trivial.
Let $\ell$ be a place of $F_{0}$. Away from finitely many bad places, there are four cases depending on its splitting behaviours in $F$ and $E_{0}$. The pairs $(\mathbf{G},\mathbf{H})$ have the following descriptions after localizing at $\ell$:
  1. (i)
    Splits in $F$ and $E_{0}$: $(\mathrm{GL}_{2n,F_{0}},\mathrm{GL}_{n,F_{0}}\times\mathrm{GL}_{n,F_{0}})$,
  2. (ii)
    Inert in $F$ and $E_{0}$: $(\mathrm{U}_{2n,F/F_{0}},\operatorname{Res}_{F/F_{0}}\mathrm{GL}_{n,F})$.
  3. (iii)
    Splits $F$, inert in $E_{0}$: $(\mathrm{GL}_{2n,F_{0}},\mathrm{GL}_{n,E_{0}})$,
  4. (iv)
    Inert in $F$, split in $E_{0}$: $(\mathrm{U}_{2n,F/F_{0}},\mathrm{GL}_{n,F_{0}}\times\mathrm{GL}_{n,F_{0}})$.
For a JNS Euler system, we need the place $\ell$ to be split in $F^{\prime}$, which corresponds to cases (i) and (ii) above. For arithmetic applications without conditions on $p$, it is necessary to consider 100% of such primes. Case (i) was handled in [GS23] using the zeta integral method. However, in case (ii), the local multiplicity is $2^{n}$. So the zeta integral method cannot be applied directly. Moreover, we are not aware of any theory of local zeta integrals for this case.
Instead, we reduce the problem to a question of unramified local harmonic analysis on the symmetric space $\mathbf{H}\backslash\mathbf{G}$, which has been studied extensively. In case (i), we use the general result of Sakellaridis [Sak13]. In case (ii), we compute the spectral decomposition ourselves following the Casselman–Shalika–Hironaka method developed in [Hir10, Off04] (Theorem 3.15). In both cases, the denominator of the Plancherel formula contains the centre of the standard $L$-function for $\mathrm{BC}_{F_{0}}^{F}(\pi)$. This is the $L$-function attached to $\mathbf{X}$ in Theorem 1.1.
Remark 1.4.
Even though we still need to do local calculations, our approach has the following advantages compared to the earlier zeta integral approach, even in cases where it is available:
  1. (1)
    The harmonic analysis of $C_{c}^{\infty}(X,\mathbb{C})$, where $X$ is a symmetric variety, is of independent interest.
  2. (2)
    In previous works, choosing the test vector and computing the zeta integral both require substantial trial and error, whereas the spectral decomposition of $C_{c}^{\infty}(X,\mathbb{C})^{K}$ in many cases can be systematically computed using the Casselman–Shalika method.
1.3. Idea of proof
Our main result is a combination of the following two steps which are completely different in nature:
  1. (1)
    Construction of a “motivic theta series” (Definition 4.1).
  2. (2)
    Local harmonic analysis on spherical varieties (Proposition 3.18).
We now explain each item in turn.
1.3.1. Motivic theta series
Observation 1.5 (Propositions 4.8, 4.11).
where $\mathbf{X}$ is a spherical variety.
More precisely, in the cycles case, we take $\mathbf{X}=\mathbf{H}\backslash\mathbf{G}$, and there is the commutative diagram
$C_{c}^{\infty}(\mathbf{G},\mathbb{Q}_{p})_{\mathbf{H}}$$\mathrm{H}_{\mathrm{cont}}^{2d}(\mathrm{Sh}_{\mathbf{G}},\mathbb{Q}_{p}(d))$$C_{c}^{\infty}(\mathbf{X},\mathbb{Z}_{p})$$\mathrm{H}_{\mathrm{cont}}^{2d}(\mathrm{Sh}_{\mathbf{G}},\mathbb{Z}_{p}(d))$}
The previous constructions worked with the top arrow directly, but the pre-image of the natural lattice in cohomology did not have an easy description. Our observation is that the bottom arrow exists (Proposition 4.8), making $C_{c}^{\infty}(\mathbf{X},\mathbb{Z}_{p})$ the natural integral lattice in the coinvariant space. The cases involving Eisenstein classes are treated similarly (Proposition 4.11). In these cases, we take $\mathbf{X}$ to be the vector bundle
This observation was already used by the third named author in [LS24] to bypass some Iwasawa-theoretic arguments needed previously to deal with torsions. This was also independently observed by A. Groutides, who used it to study the optimality of Hecke operators appearing in tame norm relations [Gro24].
1.3.2. Local harmonic analysis
Once such a motivic theta series exists, the problem of tame norm relations is entirely reduced to constructing test vectors satisfying norm relations in $C_{c}^{\infty}(X,\mathbb{Z}_{p})$, where $X=\mathbf{X}(F_{\ell})$.
(†) $\text{is it true that }\mathcal{H}_{\ell}\cdot\Phi_{0}\in\operatorname{Tr}_{K^{0}}^{K^{1}}C_{c}^{\infty}(X,\mathbb{Z}_{p})^{K^{1}}?$
This is stronger than just requiring the function $\mathcal{H}_{\ell}\cdot\Phi_{0}$ to take values in $\mathbb{Z}_{p}$. Indeed, in the extreme case where the $K^{0}$-orbit of a point $x\in X$ coincides with its $K^{1}$-orbit, we need the value at $x$ to be divisible by the index $K^{0}:K^{1}]=\ell-1$. We give a necessary condition for when this kind of additional divisibility requirements can occur, in terms of the geometry of $\mathbf{X}$ (Proposition 3.4). This is done by extending the proof of the generalized Cartan decomposition due to Gaitsgory–Nadler [GN10, Theorem 8.2.9] and Sakellaridis [Sak12, Theorem 2.3.8],
Observation 1.6 (cf. Proposition 3.11).
[…]
$\hat{\mathcal{L}}_{(X)}=(1-\ell^{-\frac{1}{2}}x)(1-\ell^{-\frac{1}{2}}x^{-1}),$
corresponding to the centre of the standard degree 2 $L$-function. The right hand side of Theorem 3.6 is
$(1-\ell^{-\frac{1}{2}}x)(1-\ell^{-\frac{1}{2}}x^{-1})\cdot\frac{1-x^{-2}}{(1-\ell^{-1/2}x^{-1})^{2}}=(1-x^{-2})\cdot\frac{1-\ell^{-1/2}x}{1-\ell^{-1/2}x^{-1}}.$
This can be expanded as a formal Laurent series in $x$. The conclusion of Theorem 3.6 is that the function $\phi\in C_{c}^{\infty}(\mathbf{X}(F)/\mathbf{G}(\mathcal{O}),\mathbb{Z}_{p})=C_{c}(\mathbb{Z}_{\leq 0},\mathbb{Z}_{p})$ whose relative Satake transform is $\hat{\mathcal{L}}_{(X)}$ is exactly given by the non-positive coefficients of this formal Laurent series, up to an explicit half power of $\ell$. By keeping track of these half powers carefully, it is easy to show that the values of $\phi$ are all polynomials in $\ell$.
From the geometric volume consideration above, we need to show that $\phi(0)\in(\ell-1)\mathbb{Z}_{p}$. To do this, we view $\ell$ as a formal variable and specialize at $\ell=1$. The rational function becomes the polynomial
$(1-x^{-2})\cdot\frac{1-x}{1-x^{-1}}=x^{-1}-x$
The coefficient of $x^{0}$ is clearly 0, which proves the desired divisibility by $\ell-1$. In general, instead of expanding out this polynomial directly, we observe that it is anti-symmetric under the reflection $x\leftrightarrow x^{-1}$. This is the strategy used to prove the general Proposition 3.11.
1.4. Arithmetic applications
Starting from Theorem 1.1, there are well-known methods in the literature to extract arithmetic consequences from it. We state one such example application, but it is not the focus of our paper.
The next corollary constructs the tame part of a JNS Euler system in Galois cohomology, following a standard Abel–Jacobi map procedure described in §4.3.
@@ 1.4. Arithmetic applications
Let $\pi$ be a stable cohomological automorphic representation of $\mathbf{G}(\mathbb{A}_{F})$ distinguished by $\mathbf{X}$. Let $\rho_{\pi}$ be the $p$-adic Galois representation attached to $\pi$ and the Shimura cocharacter for $\mathrm{Sh}_{\mathbf{G}}$.
@@
  • Kottwitz’s conjecture is used to show that $\rho_{\pi}$ actually contributes to the cohomology of $\mathrm{Sh}_{\mathbf{G}/\bar{E}}$. In all of our cases, the Shimura variety is of abelian type, and what we need is follows from the work of Kisin–Shin–Zhu [KSZ21].
@@
    In the twisted Friedberg–Jacquet case, the JNS Euler system argument does not apply directly to such decomposable representations, but we hope a suitable modification of it could yield applications towards rank 1 cases of the Bloch–Kato conjecture for $\rho_{\pi}$.
@@
    In the Eisenstein class cases, our construction gives classes over cyclotomic extensions (instead of anticyclotomic ones). Our resulting Euler system is the type considered by Rubin, cf. [LSZ22b, Definition 10.6.1]. Since we are not considering wild norm relations, the Iwasawa cohomology in loc. cit. should be replaced by the usual cohomology.